OpenStudy (anonymous):

Critical numbers of 2cosx+sin^2(x)

OpenStudy (anonymous):

\[2(-sinx)+cosx(0) + 2sinxcosx\]

OpenStudy (anonymous):

\[-2sinx+2sinxcosx=0\]

OpenStudy (anonymous):

\[cosx=\frac{ 2sinx }{ 2sinx }\]

OpenStudy (anonymous):

\[cosx=1\]

OpenStudy (anonymous):

or should i factor out 2sinx?

OpenStudy (anonymous):

\[2sinx(-1+cosx)=0\]

OpenStudy (anonymous):

sinx=0 cosx=1

OpenStudy (anonymous):

i think those are the answers. it says "critical numbers" so the answers should be >1

OpenStudy (anonymous):

the book answer says n(pi)

OpenStudy (helder_edwin):

u didn't solve \[ \large \sin x=0 \] and \[ \large \cos x=1 \]

OpenStudy (anonymous):

sin is 0 at pi and 2pi

OpenStudy (anonymous):

cos is 1 at 2pi

OpenStudy (anonymous):

not sure how they are doing this :(

OpenStudy (helder_edwin):

sin is also zero at pi

OpenStudy (anonymous):

ya so i pick what they have in common?

OpenStudy (anonymous):

which is pi?

OpenStudy (helder_edwin):

in general sin x=0 when \(x=n*\pi\)

OpenStudy (anonymous):

well cos is -1 at pi

OpenStudy (helder_edwin):

u don't have to satisfy both equations at the same time.

OpenStudy (anonymous):

hmm, so, pi(n) n is an integer, how would i explain that?

OpenStudy (helder_edwin):

when u solve \[ \large \sin x(-1+\cos x)=0 \] u have \[ \large \sin x=0\qquad\text{or}\qquad \cos x=1 \] so \[ \large x=n\pi\qquad\text{or}\qquad x=2n\pi \] ok?

OpenStudy (anonymous):

ahhh thanks @helder_edwin !!! =)

OpenStudy (helder_edwin):

whenever \(\cos x=1\), \(\sin x=0\). so the solutions of the former are included in the solutions of the later. that's why your book says \(x=n\pi\) and nothing else.

OpenStudy (helder_edwin):

u r very welcome.